OpenStudy (anonymous):

Help a gal out please? fan and Medal Charlie has a piece of toast that has cream cheese on one side, and he dropped it once. One time, it landed with the cream cheese side up. If he drops it two more times, what is the probability that it will have landed cream cheese side up a total of three times?

OpenStudy (anonymous):

(1/2)x(1/2)x(1/2) = 1/8 I believe

OpenStudy (anonymous):

oooooohh, I get it. so how much times it asks, is how much times you multiple the number given?

OpenStudy (gorv):

probability of geeting cheese side=1/2 as their are two face

OpenStudy (gorv):

each time probab=1/2 so dropped 3 times =1/2*1/2*1/2=1/8

OpenStudy (anonymous):

thanks you guys (: